OpenStudy (aferguson3):

A box of ground nutmeg weighs 1 1/3 ounce. If there are 20 teaspoons of nutmeg in the box, how much does one teaspoon of nutmeg weigh

OpenStudy (holsteremission):

\[\left(1\frac{1}{3}\frac{\text{ounce}}{\cancel{\text{box}}}\right)\times\left(\frac{1\cancel{\text{ box}}}{20\text{ teaspoons}}\right)=\frac{1\frac{1}{3}}{20}\frac{\text{ounces}}{\text{teaspoon}}=\cdots\]